RANDOM WALKS OF CIRCLE PACKINGS

Tomasz Dubejko

Abstract. A notion of random walks for circle packings is introduced. The geometry behind this notion is discussed, together with some applications. In particular, we obtain a short proof of a result regarding the type problem for circle packings, which shows that the type of a circle packing is closely related to the type of its tangency graph.

Typeset by AMS-TEX